# Benchmarking Legal Reasoning on 340 Law Exams

### v3

@@ 4.3 LLM Settings
We evaluate three representative LLM categories: (1) Reasoning LLMs include DeepSeek-R1 [13], Gemini-2.5-Pro [72], Claude-3.7-Sonnet [3], O3-mini [59], and QwQ-32B [63]; (2) Large or closed-sourced LLMs include GPT-4.1 [58], GPT-4o [55], DeepSeek-v3 [14], Llama-4-Maverick [46], GPT-4o-mini [56], and Llama-3-Instruct with 70B and 405B parameters [24]; and (3) Small and open-sourced LLMs include Gemma-3-12B-it [74], Gemma-2-9B-it [73], Phi-4 [1], EuroLLM-9B-Instruct [42], Qwen2.5-7B-Instruct [62], and Llama-3.1-8B-Instruct [47]. We select well-performing flagship LLMs from diverse families (e.g., Llama, GPT) and also prioritize those with strong multilingual abilities (e.g., EuroLLM, Gemma-3) due to LEXam’s multilingual nature.
@@
For conventional LLMs, we set temperature to 0 and max length to 4096, which is sufficient for complete answers. We adopt the lighteval framework [23] to standardize the evaluation of conventional LLMs with different endpoints. Reasoning models have more diversified sets of hyperparameters and are not supported by lighteval at the time of writing, so we follow the official recommended settings. We document detailed settings in Appendix F.
@@ 5.1 Results on Open Questions
We show our main results in Table 2. SOTA reasoning models, such as Gemini-2.5-Pro and Claude-3.7-Sonnet, achieve the best performance, with Gemini-2.5-Pro showing the highest average score (82.2). Other reasoning models, namely DeepSeek-R1, o3-mini, and QwQ-32B, are also competitive among all the models, highlighting the strength of models explicitly optimized for reasoning tasks. Among the large conventional LLMs, GPT-4.1 and GPT-4o perform strongly, with 68.2 and 66.2, respectively, followed by DeepSeek-V3 and Llama-4-Maverick. Smaller or older-generation LLMs, including reasoning models such as QwQ-32B and o3-mini, as well as conventional models like GPT-4o-mini, Llama-3.1-Instruct-405B, and Llama-3.3-Instruct-70B, generally perform worse than flagship LLMs. Small LLMs generally exhibit the worst overall performance. However, Gemma-3-12B-it stands out with an average score of 50.9, which surpasses Llama-3.1-405B-Instruct with a 33$\times$ smaller size, possibly due to the model’s particular focus on multilinguality.
Table 2: Performance of models on long-form open questions and MCQs with bootstrapped standard error. Open questions are judged by GPT-4o with temperature 0. Judge S.E. indicates the average Std of 3 samples from the GPT-4o Judge (1 with greedy decoding, 2 with temperature 0.5). Results are sorted by Score in descending order.
@@
Score + S.E. Judge S.E. Acc. + S.E.
Reasoning Gemini-2.5-Pro [72] 82.2 (± 0.35) 3.35 55.7 (± 1.18)
Claude-3.7-Sonnet [3] 77.6 (± 0.38) 3.63 57.2 (± 1.21)
DeepSeek-R1 [13] 68.4 (± 0.49) 4.16 52.4 (± 1.22)
o3-mini [59] 55.3 (± 0.43) 5.57 44.2 (± 1.23)
QwQ-32B [63] 52.9 (± 0.57) 4.21 47.8 (± 1.23)
Large GPT-4.1 [58] 68.2 (± 0.40) 5.32 54.4 (± 1.26)
GPT-4o [55] 66.2 (± 0.39) 5.56 53.1 (± 1.20)
DeepSeek-V3 [14] 60.0 (± 0.43) 5.99 46.6 (± 1.28)
Llama-4-Maverick [46] 55.4 (± 0.39) 6.14 49.1 (± 1.24)
Llama-3.1-405B-Instruct [47] 48.9 (± 0.39) 5.76 43.2 (± 1.19)
Llama-3.3-70B-Instruct [47] 46.9 (± 0.37) 5.82 28.2 (± 1.10)
Small Gemma-3-12B-it [74] 50.9 (± 0.51) 4.26 29.9 (± 1.10)
GPT-4o-mini [56] 48.6 (± 0.38) 5.80 41.0 (± 1.21)
Phi-4 [1] 43.5 (± 0.40) 5.86 40.7 (± 1.19)
Gemma-2-9B-it [73] 30.8 (± 0.37) 4.46 25.4 (± 1.04)
EuroLLM-9B-Instruct [42] 25.7 (± 0.36) 4.12 23.3 (± 1.02)
Qwen2.5-7B-Instruct [62] 18.7 (± 0.29) 4.23 29.3 (± 1.10)
Ministral-8B-Instruct [48] 16.9 (± 0.34) 4.10 26.3 (± 1.12)
Llama-3.1-8B-Instruct [47] 11.4 (± 0.27) 3.60 24.0 (± 1.05)
Across all groups, bootstrapped standard errors remain low, suggesting that the LLMs perform consistently across different subsets of the dataset. In most cases, the Judge S.E. is below 5, showing GPT-4o Judge scores are stable (on a 0–100 scale) even when sampling with a non-zero temperature.
Figure 3 shows the average LLM-Judge scores across three dimensions: Language, Legal Area, and Jurisdiction. Additional features appear in Appendix  C.1 (Figure  8). Reasoning models consistently outperform large and small models across all dimensions. For Language, all groups perform better on English than German courses, with reasoning models showing the largest gap. In Legal Area, interdisciplinary and public law topics score higher than criminal and private law. For Jurisdiction, generic and international law tasks achieve higher accuracy than Switzerland-specific ones.
Figure 3: Model performance on open questions grouped by language, legal area, and jurisdiction.
@@ 5.2 Results on Multiple-Choice Questions
Reasoning models lead in accuracy, with Claude-3.7-Sonnet (57.2%) and Gemini-2.5-Pro (55.7%) achieving the top two scores. These are closely followed by the strongest large model, GPT-4.1 (54.4%), which also demonstrates competitive performance. Among small models, GPT-4o-mini (41.0%) and Phi-4 (40.7%) outperform the rest, while most others in this category, including Gemma-2-9B-it, EuroLLM-9B-Instruct, and Llama-3.1-8B-Instruct, score below 30%. Across all groups, bootstrap standard errors are low (typically below 0.013), indicating stable performance estimates.
@@
Figure 4: Model performance on MCQs grouped by language, legal area, jurisdiction, number of statements, and question polarity.
@@ 5.2.1 Robustness Check on Multiple-Choice Questions with Perturbations
Gemini-2.5-Pro [72] 68.6 (± 2.37) 51.6 (± 2.55) 45.2 (± 2.43) 35.6 (± 2.30)
Claude-3.7-Sonnet [3] 60.9 (± 2.48) 48.6 (± 2.44) 40.4 (± 2.43) 33.0 (± 2.31)
DeepSeek-R1 [13] 57.5 (± 2.49) 44.1 (± 2.45) 36.9 (± 2.36) 24.9 (± 2.21)
GPT-4.1 [58] 58.0 (± 2.53) 42.3 (± 2.51) 33.2 (± 2.38) 26.3 (± 2.13)
GPT-4o [55] 53.7 (± 2.56) 36.4 (± 2.37) 22.6 (± 2.09) 21.8 (± 2.18)
DeepSeek-V3 [14] 58.6 (± 2.57) 36.1 (± 2.40) 28.9 (± 2.28) 16.0 (± 1.88)
o3-mini [59] 50.0 (± 2.55) 33.5 (± 2.34) 24.5 (± 2.16) 17.0 (± 1.89)
@@ 6 Expert Evaluation for Open Questions
Evaluating open-ended legal questions presents a complex challenge. Prior work has increasingly turned to LLM-based evaluators (“LLM-as-a-judge”), yet concerns remain about their reliability in high-stakes domains such as law. To address this, we introduce an LLM judge (see Section 4.1) based on GPT-4o. To rigorously validate its use, we conduct an expert evaluation. This analysis not only benchmarks the agreement between human annotators, but also assesses whether our LLM judge can serve as a statistically grounded proxy for expert annotation using the Alternative Annotator Test [8].
@@ 6.1 Inter-Annotator Agreement
We first measured how consistently our three human experts rated the same 50 randomly drawn question-answer pairs on a scale of 0-10. The annotations were done by three human evaluators, each holding a Swiss Master’s degree in law and pursuing or holding a doctorate. Two studied parts of their studies at the University of Zurich and two have passed the Zurich bar examination. They scored all 50 items blindly and independently, following only very brief calibration instructions.
@@ 6.2 Alternative Annotator Test
To rigorously assess whether our GPT-4o judge can replace human experts, we applied the alt-test [8] with $\varepsilon=0$ (i.e. no bonus). Following their guidelines, we used a leave-one-out design with $m=3$ experts and $n=50$ items. We reused the exact same 50-item random sample and three expert annotators described in Section 6.1.
In each of three leave-one-out trials, we compared GPT-4o’s distance to the mean of the other two experts against that expert’s distance, counted “wins” for the LLM versus the human, formed a difference vector $d_{i}\in\{-1,+1\}$, and ran a one-sided t-test on $\bar{d}<\varepsilon$. We then corrected the three raw $p$-values via Benjamini–Yekutieli (BY) at $\alpha=0.05$.
@@
Legal Annotator 1 33 17 0.0110 0.0303 0.66
Legal Annotator 2 29 21 0.1310 0.2402 0.58
Legal Annotator 3 37 13 0.0002 0.0010 0.74
Winning Rate $\omega$ 0.67
@@
Our GPT-4o judge surpasses the recommended threshold of $\omega\geq 0.5$. As Table 4 shows, it significantly matches or exceeds two of the three experts (Legal Annotator 1 and 3), yielding a winning rate $\omega=0.67$. Calderon et al. [8] further allow a replacement bonus $\varepsilon=0.15$ to reflect the cost savings of LLM annotation; applying this adjustment, GPT-4o even passes all three tests ($\omega=1.00$). Additional alt-test results using other LLMs are provided in Section G.3.
@@
Taken together, our expert evaluation establishes that annotators agree substantially among themselves (Pearson $r=0.70$, quadratic weighted $\kappa=0.49$, $MAE=1.95$) and that the GPT-4o judge matches ($\omega=0.67$) two of three annotators even under the strictest no-bonus criterion ($\varepsilon=0$). These results validate the use of our GPT-4o judge as a statistically justified replacement for open questions.
@@ 7 Conclusion
We introduced LEXam, a comprehensive legal reasoning benchmark composed of 340 law exams, specifically designed to challenge the performance of LLMs through both outcome- and process-based evaluations. Our analysis reveals substantial variability and limitations in LLM capabilities for addressing MCQs and especially on complex open questions; notably, increasing the number of MCQ options consistently reduces model accuracy. Our evaluation framework offers a scalable approach for assessing legal reasoning quality beyond simple accuracy metrics, thereby facilitating future research aimed at enhancing the reliability and robustness of LLMs on challenging legal tasks.
8 Limitations
The LEXam benchmark constitutes a significant step towards evaluating long-context and process-based legal reasoning. Nevertheless, several important limitations should be acknowledged. LEXam features legal questions covering numerous subjects from a wide range of law exams in both English and German. However, extending the multilingual scope further would be beneficial, particularly by incorporating questions from diverse jurisdictions representing both common law and civil law traditions. Such an expansion could enable deeper investigations into cross-legal-system and cross-cultural differences in legal reasoning performance among LLMs, which we leave for future work. Moreover, the current data curation approach relies on existing courses and examination data. Given typical timelines, legal interpretations or even the applicable laws themselves may have changed since the original course materials or questions were formulated, potentially rendering some answers outdated or incorrect over time. Fortunately, our detailed, fine-grained metadata allows easy identification of specific subsets of data impacted by subsequent legal changes. Consequently, developing a detect-and-update mechanism for continuously updating the benchmark—similar to that described by LiveLongBench [81]—is a promising future direction.
9 Acknowledgments
Thanks to Viola Meier and Julien Gaumez for conducting the initial data collection.

### v4

@@ 3.3 LLM Settings
We evaluate 26 LLMs across three representative categories: (1) Reasoning LLMs include DeepSeek-R1 (DeepSeek-AI, 2025a), GPT-5, -mini, and -nano (OpenAI, 2025b), Gemini-2.5-Pro (Gemini, 2025), Claude-3.7-Sonnet (Anthropic, 2025), O3-mini (OpenAI, 2025c), QwQ-32B (Qwen, 2024), Qwen3 32B and 235B (Qwen, 2025); (2) Large models include GPT-4.1 (OpenAI, 2025a), GPT-4o (OpenAI, 2024a), DeepSeek-V3 (DeepSeek-AI, 2025b), Llama-4-Maverick (Meta AI, 2025), and Llama-3-Instruct(it) with 70B and 405B parameters (Grattafiori et al., 2024); and (3) Small models include GPT-4.1-mini, and -nano (OpenAI, 2025a), GPT-4o-mini (OpenAI, 2024b), Gemma-3-12B-it (Gemma, 2025), Gemma-2-9B-it (Gemma, 2024), Phi-4 (Abdin et al., 2024), EuroLLM-9B-it (Martins et al., 2024), Qwen-2.5-7B-it (Qwen, 2025), Ministral-8B-it (Mistral AI, 2024), and Llama-3.1-8B-it (Meta-llama, 2024). We select well-performing flagship LLMs from diverse families (e.g., Llama, GPT) and also prioritize those with strong multilingual abilities (e.g., EuroLLM, Gemma-3) due to LEXam’s multilingual nature.
@@
For conventional LLMs, we set temperature to 0 and max length to 4096, which is sufficient for complete answers. We adopt the lighteval framework (Fourrier et al., 2023) to standardize the evaluation of conventional LLMs with different endpoints. Reasoning models have more diversified sets of hyperparameters and are not supported by lighteval at the time of writing, so we follow the official recommended settings. We document detailed settings in Appendix˜F.
@@ 4.1 Results on Open Questions
Table 1: Performance of models on long-form open questions and MCQs with bootstrapped standard error. Open questions are graded by ensembled LLM judges. Results are sorted by Judge Score in descending order.
@@
Judge Score (± S.E) Accuracy (± S.E)
Reasoning
GPT-5 70.20 (± 0.41) 62.65 (± 1.17)
Gemini-2.5-Pro 67.40 (± 0.51) 55.72 (± 1.18)
Claude-3.7-Sonnet 62.86 (± 0.51) 57.23 (± 1.21)
GPT-5-mini 60.32 (± 0.45) 54.82 (± 1.19)
DeepSeek-R1 55.91 (± 0.51) 52.41 (± 1.22)
o3-mini 48.13 (± 0.49) 44.22 (± 1.23)
Qwen3-235B 47.25 (± 0.46) 48.19 (± 1.20)
QwQ-32B 44.36 (± 0.53) 47.83 (± 1.23)
Qwen3-32B 40.00 (± 0.43) 45.30 (± 1.23)
GPT-5-nano 27.25 (± 0.63) 47.11 (± 1.19)
Large
GPT-4.1 57.50 (± 0.51) 54.40 (± 1.26)
GPT-4o 56.93 (± 0.48) 53.13 (± 1.20)
DeepSeek-V3 52.53 (± 0.48) 46.57 (± 1.28)
Llama-4-Maverick 47.25 (± 0.46) 49.10 (± 1.24)
Llama-3.1-405B-it 43.14 (± 0.41) 43.19 (± 1.19)
Llama-3.3-70B-it 41.27 (± 0.41) 28.19 (± 1.10)
Small
GPT-4.1-mini 54.58 (± 0.43) 48.49 (± 1.22)
GPT-4.1-nano 43.68 (± 0.41) 39.22 (± 1.22)
GPT-4o-mini 42.55 (± 0.39) 40.96 (± 1.21)
Gemma-3-12B-it 41.29 (± 0.48) 29.94 (± 1.10)
Phi-4 38.54 (± 0.42) 40.66 (± 1.19)
Gemma-2-9B-it 27.41 (± 0.37) 25.36 (± 1.04)
EuroLLM-9B-it 22.95 (± 0.35) 23.31 (± 1.02)
Qwen-2.5-7B-it 16.67 (± 0.29) 29.28 (± 1.10)
Ministral-8B-it 14.88 (± 0.32) 26.27 (± 1.12)
Llama-3.1-8B-it 10.00 (± 0.26) 24.04 (± 1.05)
[…]
Figure 4: Model performance on open questions grouped by metadata.
Figure 4 shows average LLM-Judge scores across three dimensions: Language, Legal Area, and Jurisdiction. Additional features appear in Appendix  C.1 (Figure  9). Reasoning models consistently outperform large and small models universally. For Language, all groups perform better on English than German courses, with small models showing the largest gap. In Legal Area, interdisciplinary and public law topics score higher than criminal and private law. For Jurisdiction, generic and international law tasks achieve higher accuracy than Switzerland-specific ones.
@@ 4.2 Results on Multiple-Choice Questions
Reasoning models lead in accuracy, with GPT-5 (62.65%) and Claude-3.7-Sonnet (57.23%) achieving the top two scores. These are closely followed by the strongest large model, GPT-4.1 (54.40%), which also demonstrates competitive performance. Among small models, GPT-4.1-mini (48.49%) and GPT-4o-mini (40.96%) outperform the rest, while most others in this category, including Gemma-2-9B-it, EuroLLM-9B-it, and Llama-3.1-8B-it, score below 30%. Across all groups, bootstrap standard errors are low (typically below 1.3%), indicating stable performance estimates.
@@
Figure 5: Model performance on MCQs grouped by various metadata.
@@ 4.2.1 Robustness Check on Multiple-Choice Questions with Perturbations
Gemini-2.5-Pro 68.61 (± 2.37) 51.56 (± 2.55) 45.24 (± 2.43) 35.62 (± 2.30)
Claude-3.7-Sonnet 60.92 (± 2.48) 48.59 (± 2.44) 40.38 (± 2.43) 33.02 (± 2.31)
DeepSeek-R1 57.54 (± 2.49) 44.11 (± 2.45) 36.94 (± 2.36) 24.93 (± 2.21)
GPT-4.1 58.02 (± 2.53) 42.31 (± 2.51) 33.23 (± 2.38) 26.30 (± 2.13)
GPT-4o 53.73 (± 2.56) 36.42 (± 2.37) 22.55 (± 2.09) 21.81 (± 2.18)
DeepSeek-V3 58.57 (± 2.57) 36.07 (± 2.40) 28.92 (± 2.28) 16.03 (± 1.88)
o3-mini 50.02 (± 2.55) 33.54 (± 2.34) 24.46 (± 2.16) 17.01 (± 1.89)
@@ 5 Expert Evaluation for Open Questions
Evaluating open-ended legal questions presents a complex challenge. Prior work has increasingly turned to LLM-based evaluators (“LLM-as-a-judge”), yet concerns remain about their reliability in high-stakes domains such as law. To address this, we introduce an ensemble LLM judge (see Section˜3.1 and Appendix G.4 for more details) based on GPT-4o, Qwen3-32B, and DeepSeek-V3. To rigorously validate its use, we conduct an expert evaluation. This analysis both benchmarks human annotator agreement and assesses whether our LLM judge can serve as a statistically grounded proxy for expert annotation using the Alt-test (Calderon et al., 2025).
@@ 5.1 Inter-Annotator Agreement
We first measured the consistency of three legal experts in rating 50 randomly drawn question–answer pairs on a 0-10 scale. All three hold a Swiss Master’s degree in law and are pursuing or have completed a doctorate in Law. The experts scored all 50 items blindly and independently, following only brief calibration instructions (see Appendix G.1 for details).
@@ 5.2 Alternative Annotator Test
To rigorously assess whether our ensemble LLM judge can replace human experts, we applied the Alt-test (Calderon et al., 2025) with $\varepsilon{=}0$ (i.e. no bonus or tolerance margin for LLM judges). Following their guidelines, we used a leave‑one‑out design with $m=3$ experts and n=50n{=}50 items. We reused the exact same 50‑item random sample and three expert annotators described in Section˜5.1.
In each of three leave‑one‑out trials, we compared our ensemble LLM judge’s distance to the mean of the other two experts against that expert’s distance, counted “wins” for the LLM versus the human, formed a difference vector $d_{i}\in\{-1,+1\}$, and ran a one‑sided t‑test on $\bar{d}<\varepsilon$. We then corrected the three raw $p$-values via Benjamini–Yekutieli (BY) at $\alpha{=}0.05$.
@@
Legal Expert 1 36 14 0.001 0.002 0.72
Legal Expert 2 34 16 0.005 0.009 0.68
Legal Expert 3 37 13 0.000 0.001 0.74
Winning Rate $\omega$ 1.00
@@
Our ensemble LLM judge surpasses the recommended threshold of $\omega\geq 0.5$. As Table 3 shows, it significantly exceeds the three experts (Legal Annotator 1 and 3), yielding a winning rate $\omega{=}1.00$. Additional Alt-test results using other LLMs are provided in Section˜G.4.
@@
Taken together, our expert evaluation establishes that annotators agree substantially among themselves (Pearson r=0.70r{=}0.70, quadratic weighted $\kappa{=}0.49$, MAE=1.95MAE{=}1.95) and that our ensemble LLM judge matches ($\omega{=}1.00$) three annotators even under the strictest no-bonus criterion ($\varepsilon{=}0$). These results validate our ensemble LLM judge as a statistically justified replacement for open questions.
5.3 Further Robustness Analyses of LLM Judges
We conducted additional quantitative and qualitative analyses to assess the robustness of judge models across languages and legal domains using four models (GPT-4o, Gemini-2.5-Pro, DeepSeek-R1, and Claude-4-Sonnet). Correlation tests on expert-annotated samples show that all models align closely with human scores in both English and German, with only minor and statistically insignificant language effects, indicating cross-linguistic consistency (see Tables 14 and 15). Qualitative inspection of LLM answers revealed recurring failure modes, including flawed legal reasoning, references to incorrect or fictitious authorities, and weaker German or multilingual proficiency in smaller models, which often resulted in incoherent outputs. We also observed systematic differences in leniency: GPT-4o was the most generous, Claude-4-Sonnet and Gemini-2.5-Pro were stricter, and DeepSeek-R1 fell in between. Notably, the largest gaps between model and human evaluations appeared in cases with higher human disagreement, underscoring the inherent difficulty of borderline legal judgments. These results further motivate the use of an ensemble judge with minimum-score integration. More details are available in Appendices H and I.
6 Related Work
6.1 International Benchmarks
Multiple international benchmarks for legal reasoning have been established, comprising various individual datasets, primarily sourced from China and the United States. Examples are LegalBench (Guha et al., 2023), LawBench (Fei et al., 2023), LexEval (Li et al., 2024), LEXTREME (Niklaus et al., 2023b) or LexGLUE (Chalkidis et al., 2022). While they propel the field forward, they still fall short in the complexity of the tasks evaluated. In contrast, we introduce very challenging law exam questions, often requiring detailed, long-form answers.
In Table˜4, we compare LEXam to prior legal QA datasets. We demonstrate that LEXam a) covers a wide array of legal disciplines (78 subdomains), b) is the first general legal reasoning dataset containing challenging long-form ground-truth answers, and c) is the only dataset with an expert-tuned LLM-as-a-judge system for evaluation.
6.2 Swiss Benchmarks
Research on NLP benchmarks around Swiss legal data started with the task of legal judgment prediction, where models predict the case outcome based on facts (Niklaus et al., 2021, 2022). Subsequently, many other tasks have been studied on Swiss court rulings, such as explainability for judgment prediction (T.y.s.s. et al., 2024), negation scope resolution Christen et al. (2024), sentence boundary detection (Brugger et al., 2023), citation extraction (Rasiah et al., 2023), case importance prediction (Stern et al., 2024), law area prediction (Rasiah et al., 2023), summarization of leading decisions (Rolshoven et al., 2024), drafting of considerations (Rasiah et al., 2023), translation (Niklaus et al., 2025), and the analysis of gender distribution in court decisions (Merane, 2021). Additionally, prior work studied the anonymization (Niklaus et al., 2023a) and risk of re-identification of involved persons in court rulings (Nyffenegger et al., 2024).
Most prior tasks focus on content from court rulings and, to some extent, legislation. To our knowledge, no Swiss legal dataset from law school exams exists. In much prior work, labels have been extracted through regexes instead of manual human annotation (Merane & Geering, 2023). While regexes allow for very large-scale datasets, the covered tasks may not be as natural compared to fully human-annotated datasets. In this work, we collect a large dataset of real-world university law school exams, including detailed reference responses written by leading law professors.
Housing Statutes Unknown US English Statutory Housing Law 6,853 TF
Sara Unknown US English Tax Liability 160 TF
Brazilian Bar Exams Unknown Brazil Portuguese 17 law areas 2,130 MC
COLIEE CC BY 4.0 Japan, Canada Japanese, English Japanese Civil Law, Canadian Case Law 1,774 MC
GLOBALCIT CC BY 4.0 International English International Citizenship Law 9,310 MC
JecQA CC BY-NC-ND China Chinese Chinese Law 21,072 MC
MMLU (Legal Subset) CC BY 4.0 Primarily US English General Legal Knowledge 1,970 MC
Multistate Bar Exam Unknown US English Bar Exam questions 1,195 MC
PrivacyQA MIT Global English Privacy Law & Data Protection 1,750 LF (140)
@@ 7 Conclusion
We introduced LEXam, a comprehensive legal reasoning benchmark composed of 340 law exams, specifically designed to challenge the performance of LLMs through both outcome- and process-based evaluations. Our analysis reveals substantial variability and limitations in LLM capabilities for addressing MCQs and especially on complex open questions; notably, increasing the number of MCQ options consistently reduces model accuracy. Our evaluation framework offers a scalable approach for assessing legal reasoning quality beyond simple accuracy metrics, thereby facilitating future research aimed at enhancing the reliability and robustness of LLMs on challenging legal tasks. Plans for further data collection are detailed in Appendix J.
Ethics Statement
This work introduces a benchmark for evaluating LLMs on legal reasoning tasks using law school exam data. All exam questions and reference answers were obtained with permission from the University of Zurich Faculty of Law and are used exclusively for research purposes. The dataset consists solely of publicly available academic materials and excludes any personal, confidential, or sensitive content, such as student-written answers.
Our analyses focus exclusively on model performance and do not constitute legal advice. Moreover, while the benchmark reflects authentic legal reasoning tasks, it is intended purely as a research and educational resource, not as a tool for any real-world legal practice. Given the high-stakes nature of the legal domain, we strongly emphasize that current LLMs must not be relied upon for legal decision-making without qualified human oversight.
Reproducibility Statement
To facilitate reproducibility and future research, we release both the full LEXam dataset and the complete evaluation pipeline. Specifically, (a) the dataset—which includes all exam questions, metadata, and reference materials—and (b) the evaluation code are provided in the anonymous repository and uploaded as supplementary materials. Together, these resources enable replication of our results and support extensions of the benchmark to new models, tasks, and evaluation methods.